# Quantum Enhanced Sensitivity through Many-Body Bloch Oscillations

### v1

@@ Abstract
We investigate the sensing capacity of non-equilibrium dynamics in quantum systems exhibiting Bloch oscillations. By focusing on resource efficiency of the probe, quantified by quantum Fisher information, we find different scaling behaviors in two different phases, namely localized and extended. Our results provide a quantitative ansatz for quantum Fisher information in terms of time, probe size, and the number of excitations. In the long-time regime, the quantum Fisher information is a quadratic function of time, touching the Heisenberg limit. The system size scaling drastically depends on the phase changing from super-Heisenberg scaling in the extended phase to size-independent behavior in the localized phase. Furthermore, increasing the number of excitations always enhances the precision of the probe, although, in the interacting systems the enhancement becomes less eminent than the non-interacting probes, which is due to induced localization by interaction between excitations.
@@
Introduction.—The delicate nature of quantum systems makes them naturally suitable for sensing gravitational, magnetic and electric fields with unprecedented precision well beyond the capacity of classical probes [1]. The ultimate precision of a probe for sensing an unknown parameter $h$, quantified by standard deviation $\delta h$, is fundamentally bounded by Cramér-Rao inequality $\delta h{\geq}1/\sqrt{\mathcal{M}\mathcal{F}_{Q}}$, where $\mathcal{M}$ is the number of samples and $\mathcal{F}_{Q}$ is Quantum Fisher Information (QFI) [2, 3]. The equality can only be achieved when the probe is measured in an optimal basis and thus the QFI indeed dictates the ultimate achievable precision. The performance of a sensor is determined by the scaling of its QFI with respect to resources (e.g. probe size and time), namely $\mathcal{F}_{Q}{\sim}L^{\beta}$, where $L$ is the resource and $\beta$ is an exponent. While in the absence of quantum features, one can at best achieve $\beta{=}1$ (i.e. standard limit), in quantum probes, the precision might be enhanced to $\beta>1$, known as quantum enhanced sensitivity, in which a special case is $\beta{=}2$ (i.e. Heisenberg limit) [4].
@@
Quantum criticality is known to be instrumental for achieving quantum enhanced sensitivity. In fact, various forms of criticality have been exploited for sensing purposes, including first-order [5, 6, 7], second-order [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24], dissipative [25, 26, 27, 28, 29, 30, 31, 32, 33, 34], topological [35, 36, 37, 38], Floquet [39, 40], time crystals [41, 42, 43], non-Hermitian systems [44, 45, 46, 47, 48, 49], and Stark [50, 51, 52] phase transitions. However, despite a few attempts for realizing criticality-based sensors [53, 19], their experimental implementation faces real challenges. Normally, in most of such sensors the probe should be initialized in one of its eigenstates, e.g. the ground state, which is very challenging in practice and may require extra resources to be achieved [18]. In addition, the parameter interval over which such criticality-based probe achieves quantum enhanced sensitivity is often very narrow covering only around the phase transition point. A possible solution for addressing these challenges is to go beyond equilibrium physics and exploit non-equilibrium dynamics of many-body systems [54, 55, 56, 57, 58], which is easy to implement in various physical platforms [59, 60, 61, 62, 63]. Such systems do not require complex initialization and may allow quantum enhanced sensitivity over a wider range. Several open questions exist, including: (i) what types of non-equilibrium systems may lead to quantum enhanced sensitivity? and (ii) can such probes operate optimally across an entire phase of matter rather than a narrow region around the phase transition point?
@@
Bloch oscillation is a fundamental phenomenon in condensed matter physics in which a particle under the impact of a gradient potential (i.e. a constant force) oscillates in a regular lattice [64]. The gradient potential naturally makes the neighboring lattice sites off-resonant and hence suppresses the tunneling of particles. Unlike classical systems in which particles simply move in the direction of the external force, this suppression of tunneling in quantum systems induces Wannier-Stark localization in the system [65]. The localization properties of Wannier-Stark systems have been extensively studied in both theory [66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89] and experiments [90, 91, 92, 93]. Bloch oscillations have also been observed in semiconductor devices [94], optical waveguides [95, 96, 97], cold atoms in optical lattices [98] and superconducting simulators [99, 100]. Unlike the conventional second-order quantum phase transitions which only affect the ground state, the Wannier-Stark localization takes place across the entire spectrum. This implies that the impact of such transition should be observable in non-equilibrium dynamics, where many eigenstates involve. Recently, it has been shown that individual eigenstates of the Wannier-Stark systems can be used for sensing the gradient field with super-Heisenberg precision [50]. A natural question is whether one can exploit experimental-friendly non-equilibrium dynamics of such systems for sensing purposes.
@@
In this letter, we explore the sensing capacity of the Bloch oscillations in single- and many-body Stark systems. We provide a comprehensive analysis for the scaling of the QFI in terms of time, probe size, and the number of excitations. We show that indeed many-body systems with Bloch oscillations may allow for sensing precision with super Heisenberg scaling. Unlike critical probes at equilibrium we do not demand complex initialization and the probe operates optimally over the entire extended phase.
@@
Bloch oscillation in Stark systems.— We establish our theory with the simple case of a single excitation in a one-dimensional lattice of size $L$ described by the Hamiltonian ($\hbar=1$)
$H=-J\sum_{l=1}^{L-1}\mathinner{|{l}\rangle}\!\mathinner{\langle{l+1}|}+\mathinner{|{l+1}\rangle}\!\mathinner{\langle{l}|}+h\sum_{l=1}^{L}l\mathinner{|{l}\rangle}\!\mathinner{\langle{l}|}~{},$ (1)
@@
where $J$ is the exchange coupling, $h$ is the gradient field and $\mathinner{|{l}\rangle}$ represents the excitation at site $l$. The gradient field induces a position-dependent energy shift on each site which suppresses the particle tunneling. This leads to the localization of the wave function into a limited number of sites, known as Wannier-Stark localization [65]. In the limit of $L{\gg}1$, the eigenstates of Hamiltonian (1), the Wannier-Stark states, are found as $\mathinner{|{E_{m}}\rangle}=\sum_{l=m}^{L}\mathcal{J}_{l}\big{(}\frac{2J}{h}\big{)}\mathinner{|{l}\rangle}$ whose corresponding eigenvalues are $E_{m}=mh$ [66]. Here, $\mathcal{J}_{l}(\cdot)$ are the Bessel function of the first kind. The state $\mathinner{|{E_{m}}\rangle}$ is a superposition of the states centered at the $m$th site and spreading to the neighboring sites with the extend about $4J{/}h$ lattice periods [101]. Any initial state $\mathinner{|{\Psi(0)}\rangle}=\sum_{l}f_{l}\mathinner{|{l}\rangle}$, evolves under the action of the Hamiltonian $\mathinner{|{\Psi(t)}\rangle}=e^{-iHt}\mathinner{|{\Psi(0)}\rangle}$ which can be cast into $\mathinner{|{\Psi(t)}\rangle}=\sum_{l,l^{\prime}=1}^{L}\mathcal{K}_{l,l^{\prime}}(t)f_{l^{\prime}}\mathinner{|{l}\rangle}$ with the evolution propagator $\mathcal{K}_{l,l^{\prime}}(t){=}\sum_{m=1}^{L}\braket{l}{E_{m}}\!\braket{E_{m}}{l^{\prime}}e^{-iE_{m}t}$ [66]. Note that in the large $L$ limit, the energies are equidistant, i.e. $\Delta E{\equiv}E_{m+1}{-}E_{m}{=}h$. Hence, the evolution propagator exhibits a time-periodic behavior with the characteristic time $T_{\text{Bloch}}=2\pi/h$, that leads to the Bloch oscillations. In order to see the dynamical behavior of the system, we depict $P_{l}(t){=}|\langle{l}|\Psi(t)\rangle|^{2}$ as a function of time in Figs. 1(a)-(c) for three different values of $h$ in a system of size $L{=}100$, initialized at the central site. For larger values of $h$, shown in Fig. 1(a), one clearly observes periodic spreading of the excitation over a finite distance and returning to its original site. Here, the Bloch oscillations indicate a localized phase. By decreasing $h$ the extent of periodic wave packet spreading increases until $h{=}h_{c}{\simeq}8J{/}L$, at which the wave packet delocalizes across the entire system and then localizes back, see Fig. 1(b). This is the point of transition in which the localized phase transforms into an extended phase. Indeed, by further decreasing of $h$, a new behavior emerges as the excitation spreads across the entire system but does not fully return to its initial localized site, signaling an ergodic behavior [101, 50].
@@
[…]
@@
$\mathcal{F}_{Q}$ $\sim h^{-2}t^{2}L^{0},\;\;\text{ for }h>h_{c}~{},$
$\mathcal{F}_{Q}$ $\sim t^{2}L^{\beta},\hskip 21.33955pt\text{ for }h\leq h_{c}~{},$ (2)
where the scaling exponent $\beta$ may vary with $h$. Indeed, at the transition point $h_{c}{\simeq}8J/L$, the QFI from both sides of the transition point, given in Eqs. (2), match with each other which results in $\beta{=}2$. This is confirmed numerically as for different system sizes we find maximized $\mathcal{F}_{Q}/t^{2}$ with respect to $h$ at large time $t$ where $\mathcal{F}_{Q}/t^{2}$ is saturated, see the green squares in Fig. 2(b). Interestingly, in the extended phase, the scaling becomes even better as for $h/J{=}0.01$ the exponent $\beta$ increases to $\beta\simeq 2.5$, depicted by blue triangles in Fig. 2(b). Finally, to better understand the metrological aspects of our model we extract the scaling exponent $\beta$ for arbitrary values of $h$, see Supplemental Materials (SM) for more details. The results are shown in Fig. 2(c), where one can see a clear super-Heisenberg scaling with $\beta>2$ in the entire extended phase. The maximum scaling is obtained around $hL/J{\simeq}5$ whose exponent is $\beta{\simeq}3.5$. As mentioned before, at the transition point $h{=}h_{c}$ the QFI scales with Heisenberg precision, i.e. $\beta{=}2$, which then drops to $\beta{=}0$ in the localized phase.
@@
Model 2: many-body probe (half filling).—We now go beyond single excitation and extend our results to systems that consist of multi-particles which may interact with each other. We consider the XXZ Hamiltonian
$H=-\frac{J}{2}\sum_{l=1}^{L-1}(\sigma^{x}_{l}\sigma_{l+1}^{x}+\sigma^{y}_{l}\sigma_{l+1}^{y}+\Delta\sigma^{z}_{l}\sigma_{l+1}^{z})+h\sum_{l=1}^{L}l\sigma^{+}_{l}\sigma^{-}_{l}~{},$ (3)
where $\Delta$ represents the strength of the interaction between particles and $\sigma_{l}^{x,y,z}$ are the Pauli matrices acting on site $l$. In the single excitation subspace and with $\Delta{=}0$, the Hamiltonian in Eq. (3) reduces to the single-particle case in Eq. (1). We consider the time evolution of a system initially prepared in a Neel state, i.e. $\mathinner{|{\Psi(0)}\rangle}=\mathinner{|{1,0,\cdots,1,0}\rangle}$ for even $L$ and $\mathinner{|{\Psi(0)}\rangle}=\mathinner{|{1,0,\cdots,0,1}\rangle}$ for odd $L$. More details about the time evolution and complex pattern of many-body Bloch oscillations are available in SM. We first put our focus on $\Delta{=}0$ (non-interacting case) and $\Delta{=}1$ (isotropic Heisenberg Hamiltonian). The results for the more general case of $10{<}\Delta{<}$ 1 are presented and discussed in the next section. In Figs. 3(a) and 3(b), we plot the long time limit of $\mathcal{F}_{Q}/t^{2}$ as a function of $h/J$ for various system sizes $L$ for $\Delta{=}0$ and $\Delta{=}1$, respectively. In both interacting and non-interacting regimes, the normalized QFI peaks at the transition point $h_{c}$ and it shows algebraic decay of the form $\mathcal{F}_{Q}/t^{2}\sim h^{-2}$ in the localized phase, which is similar to the single excitation case. However, unlike the single excitation scenario, even in the localized phase the QFI shows size dependence. Furthermore, one can see that in the presence of interaction, the behavior of the QFI qualitatively changes as: (i) it shows fluctuations in the vicinity of the transition point; and (ii) its value becomes smaller than the non-interacting case (i.e. $\Delta=0$). To extract the scaling with system size, in Figs. 3(c) and 3(d) we plot the long-time value of $\mathcal{F}_{Q}/t^{2}$ at $h{=}h_{c}$ as a function of system size $L$ for $\Delta{=}0$ and $\Delta{=}1$, respectively. Interestingly, in both cases, the scaling suggests $\beta>2$, which indicates a super-Heisenberg scaling. Especially, in the non-interacting case $\Delta=0$, it takes the form of $\mathcal{F}_{Q}(h_{c})\sim t^{2}L^{3}$. The difference of $\beta>2$ here, compared to the single-particle result of $\beta{=}2$ at the transition point, is related to the number of excitations that will be further explained in the next section.

### v2

@@ Abstract
We investigate the sensing capacity of non-equilibrium dynamics in quantum systems exhibiting Bloch oscillations. By focusing on the resource efficiency of the probe, quantified by quantum Fisher information, we find different scaling behaviors in two different phases, namely localized and extended. Our results provide a quantitative ansatz for quantum Fisher information in terms of time, probe size, and the number of excitations. In the long-time regime, the quantum Fisher information is a quadratic function of time, touching the Heisenberg limit. The system size scaling drastically depends on the phase changing from super-Heisenberg scaling in the extended phase to size-independent behavior in the localized phase. Furthermore, increasing the number of excitations always enhances the precision of the probe, although, in the interacting systems the enhancement becomes less eminent than the non-interacting probes. This is due to the induced localization by increasing the interaction between the excitations.
1 Introduction
@@
The delicate nature of quantum systems makes them naturally suitable for sensing gravitational, magnetic, and electric fields with unprecedented precision well beyond the capacity of classical probes [1]. In order to estimate an unknown parameter $h$, encoded in the quantum state of a probe $\rho(h)$, one has to perform measurement. For a given measurement, described by a set of projectors $\{\pi_{x}\}$, the results follow a classical probability distribution in which any outcome appears with the probability $p_{x}(h)=\mathrm{Tr}[\pi_{x}\rho(h)]$. In this situation, the precision of the probe for sensing parameter $h$, quantified by standard deviation $\delta h$, is fundamentally bounded by Cramér-Rao inequality $\delta h{\geq}1/\sqrt{\mathcal{M}\mathcal{F}_{C}}$, where $\mathcal{M}$ is the number of samples and $\mathcal{F}_{C}$ is called Classical Fisher information (CFI) which is defined as [2, 3, 4]
$\mathcal{F}_{C}(h)=\sum_{x}\frac{1}{p_{x}(h)}(\frac{dp_{x}}{dh})^{2}.$ (1)
In order to obtain a measurement-independent quantity, one can maximize the CFI with respect to all possible measurements to obtain Quantum Fisher Information (QFI), namely $\mathcal{F}_{Q}=\underset{\{\pi_{x}\}}{\mathrm{max}}\mathcal{F}_{C}$. The QFI indeed describes the ultimate precision limit for which the Cramér-Rao inequality can be written as
$\delta h\geq\frac{1}{\sqrt{\mathcal{M}\mathcal{F}_{C}}}\geq\frac{1}{\sqrt{\mathcal{M}\mathcal{F}_{Q}}},$ (2)
Interestingly, the maximization over all possible measurements in the definition of the QFI can be reduced to a closed formula [3]. For instance, in the case of pure states $\rho(h){=}\mathinner{|{\Psi(h)}\rangle}\mathinner{\langle{\Psi(h)}|}$ the QFI takes the form [3]
$\mathcal{F}_{Q}(h)=4[\braket{\partial_{h}\Psi}{\partial_{h}\Psi}-|\braket{\partial_{h}\Psi}{\Psi}|^{2}].$ (3)
The performance of a sensor is determined by the scaling of its QFI with respect to resources (probe size and time), namely $\mathcal{F}_{Q}{\sim}L^{\beta}$, where $L$ is the resource and $\beta$ is an exponent. While in the absence of quantum features, one can at best achieve $\beta{=}1$ (i.e. standard limit), in quantum probes, the precision might be enhanced to $\beta>1$, known as quantum enhanced sensitivity, in which a special case is $\beta{=}2$ (i.e. Heisenberg limit) [5].
@@
Quantum criticality is known to be instrumental for achieving quantum enhanced sensitivity. In fact, various forms of criticality have been exploited for sensing purposes, including first-order [6, 7, 8], second-order [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25], dissipative [26, 27, 28, 29, 30, 31, 32, 33, 34, 35], topological [36, 37, 38, 39], Floquet [40, 41], time crystals [42, 43, 44], non-Hermitian systems [45, 46, 47, 48, 49, 50], and Stark [51, 52, 53] phase transitions. However, despite a few attempts for realizing criticality-based sensors [54, 20], their experimental implementation faces real challenges. Normally, in most of such sensors, the probe should be initialized in one of its eigenstates, e.g. the ground state, which is very challenging in practice and may require extra resources to be achieved [19]. In addition, the parameter interval over which such criticality-based probe achieves quantum enhanced sensitivity is often very narrow covering only around the phase transition point. A possible solution for addressing these challenges is to go beyond equilibrium physics and exploit non-equilibrium dynamics of many-body systems [55, 56, 57, 58, 59], which is easy to implement in various physical platforms [60, 61, 62, 63, 64]. Such systems do not require complex initialization and may allow quantum enhanced sensitivity over a wider range. Several open questions exist, including: (i) what types of non-equilibrium systems may lead to quantum enhanced sensitivity? and (ii) can such probes operate optimally across an entire phase of matter rather than a narrow region around the phase transition point?
@@
Bloch oscillation is a fundamental phenomenon in condensed matter physics in which a particle under the impact of a gradient potential (i.e. a constant force) oscillates in a regular lattice [65]. The gradient potential naturally makes the neighboring lattice sites off-resonant and hence suppresses the tunneling of particles. Unlike classical systems in which particles simply move in the direction of the external force, this suppression of tunneling in quantum systems induces Wannier-Stark localization in the system [66]. The localization properties of Wannier-Stark systems have been extensively studied in both theory [67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90] and experiments [91, 92, 93, 94]. Bloch oscillations have also been observed in semiconductor devices [95], optical waveguides [96, 97, 98], cold atoms in optical lattices [99] and superconducting simulators [100, 101]. Unlike the conventional second-order quantum phase transitions which only affect the ground state, the Wannier-Stark localization takes place across the entire spectrum. This implies that the impact of such transition should be observable in non-equilibrium dynamics, where many eigenstates are involved. Recently, it has been shown that individual eigenstates of the Wannier-Stark systems can be used for sensing the gradient field with super-Heisenberg precision [51]. A natural question is whether one can exploit experimental-friendly non-equilibrium dynamics of such systems for sensing purposes.
@@
In this paper, we explore the sensing capacity of the Bloch oscillations in single- and many-body Stark systems. We provide a comprehensive analysis for the scaling of the QFI in terms of time, probe size, and the number of excitations. We show that indeed many-body systems with Bloch oscillations may allow for sensing precision with super Heisenberg scaling. Unlike critical probes at equilibrium we do not demand complex initialization and the probe operates optimally over the entire extended phase.
2 Bloch oscillation in Stark systems
@@
We establish our theory with the simple case of a single excitation in a one-dimensional lattice of size $L$ described by the Hamiltonian ($\hbar=1$)
$H=-J\sum_{l=1}^{L-1}\mathinner{|{l}\rangle}\!\mathinner{\langle{l+1}|}+\mathinner{|{l+1}\rangle}\!\mathinner{\langle{l}|}+h\sum_{l=1}^{L}l\mathinner{|{l}\rangle}\!\mathinner{\langle{l}|}\leavevmode\nobreak\ ,$ (4)
@@
where $J$ is the exchange coupling, $h$ is the gradient field and $\mathinner{|{l}\rangle}$ represents the excitation at site $l$. The gradient field induces a position-dependent energy shift on each site which suppresses the particle tunneling. This leads to the localization of the wave function into a limited number of sites, known as Wannier-Stark localization [66]. In the limit of $L{\rightarrow}\infty$, the eigenstates of Hamiltonian (4), the Wannier-Stark states, are found as $\mathinner{|{E_{m}}\rangle}=\sum_{l=1}^{L}\mathcal{J}_{l-m}\big{(}\frac{2J}{h}\big{)}\mathinner{|{l}\rangle}$ whose corresponding eigenvalues are $E_{m}=mh$ [67]. Here, $\mathcal{J}_{l}(\cdot)$ are the Bessel function of the first kind. The state $\mathinner{|{E_{m}}\rangle}$ is a superposition of the states centered at the $m$th site and spreading to the neighboring sites with the extend about $4J{/}h$ lattice periods [102]. Any initial state $\mathinner{|{\Psi(0)}\rangle}=\sum_{l}f_{l}\mathinner{|{l}\rangle}$, evolves under the action of the Hamiltonian $\mathinner{|{\Psi(t)}\rangle}=e^{-iHt}\mathinner{|{\Psi(0)}\rangle}$ which can be cast into $\mathinner{|{\Psi(t)}\rangle}=\sum_{l,l^{\prime}=1}^{L}\mathcal{K}_{l,l^{\prime}}(t)f_{l^{\prime}}\mathinner{|{l}\rangle}$ with the evolution propagator $\mathcal{K}_{l,l^{\prime}}(t){=}\sum_{m=1}^{L}\braket{l}{E_{m}}\!\braket{E_{m}}{l^{\prime}}e^{-iE_{m}t}$ [67]. Note that in the large $L$ limit, the energies are equidistant, i.e. $\Delta E{\equiv}E_{m+1}{-}E_{m}{=}h$. Hence, the evolution propagator exhibits a time-periodic behavior with the characteristic time $T_{\text{Bloch}}=2\pi/h$, which leads to the Bloch oscillations. In order to see the dynamical behavior of the system, we depict $P_{l}(t){=}|\langle{l}|\Psi(t)\rangle|^{2}$ as a function of time in Figs. 1(a)-(c) for three different values of $h$ in a system of size $L{=}100$, initialized at the central site. For larger values of $h$, shown in Fig. 1(a), one clearly observes the periodic spreading of the excitation over a finite distance and returning to its original site. Here, the Bloch oscillations indicate a localized phase. By decreasing $h$ the extent of periodic wave packet spreading increases until $h{=}h_{c}{\simeq}8J{/}L$, at which the wave packet delocalizes across the entire system and then localizes back, see Fig. 1(b). This is the point of transition in which the localized phase transforms into an extended phase. Indeed, by further decreasing of $h$, a new behavior emerges as the excitation spreads across the entire system but does not fully return to its initial localized site, signaling an ergodic behavior [102, 51].
@@
3 Model 1: Single excitation probe
A direct consequence of the above evolution is to imprint the information of $h$ in the quantum state $\mathinner{|{\Psi(t)}\rangle}$, which is clearly evidenced in Figs. 1(a)-(c). In Fig. 1(d), we plot the normalized QFI, namely $\mathcal{F}_{Q}/t^{2}$, as a function of time for three values of $h$. One notices a rapid growth in time and eventually saturation indicating that after a transient time, the QFI eventually scales quadratically, i.e. $\mathcal{F}_{Q}{\sim}t^{2}$. Crucially, for a given length we observe that the highest long-time value of $\mathcal{F}_{Q}/t^{2}$ is attained when the system is tuned close to its transition point $h{=}h_{c}$. Indeed, this is the regime that exhibits the widest sustainable superposition of the wave packet $\mathinner{|{\Psi(t)}\rangle}$. Also, Fig 1(d) shows the size-independent scaling in early times of the QFI as $\mathcal{F}_{Q}{\sim}t^{4}$.
@@ 3 Model 1: Single excitation probe
$\mathcal{F}_{Q}$ $\sim h^{-2}t^{2}L^{0},\;\;\text{ for }h>h_{c}\leavevmode\nobreak\ ,$
$\mathcal{F}_{Q}$ $\sim t^{2}L^{\beta},\hskip 21.33955pt\text{ for }h\leq h_{c}\leavevmode\nobreak\ ,$ (5)
where the scaling exponent $\beta$ may vary with $h$. Indeed, at the transition point $h_{c}{\simeq}8J/L$, the QFI from both sides of the transition point, given in Eqs. (5), match with each other which results in $\beta{=}2$. This is confirmed numerically as for different system sizes we find maximized $\mathcal{F}_{Q}/t^{2}$ with respect to $h$ at large time $t$ where $\mathcal{F}_{Q}/t^{2}$ is saturated, see the green squares in Fig. 2(b). Interestingly, in the extended phase, the scaling becomes even better as for $h/J{=}0.01$ the exponent $\beta$ increases to $\beta\simeq 2.5$, depicted by blue triangles in Fig. 2(b). Finally, to better understand the metrological aspects of our model we extract the scaling exponent $\beta$ for arbitrary values of $h$, see Appendix A for more details. The results are shown in Fig. 2(c), where one can see a clear super-Heisenberg scaling with $\beta>2$ in the entire extended phase. The maximum scaling is obtained around $hL/J{\simeq}5$ whose exponent is $\beta{\simeq}3.5$. As mentioned before, at the transition point $h{=}h_{c}$ the QFI scales with Heisenberg precision, i.e. $\beta{=}2$, which then drops to $\beta{=}0$ in the localized phase.
@@
4 Model 2: many-body probe (half filling)
We now go beyond single excitation and extend our results to systems that consist of multi-particles which may interact with each other. We consider the XXZ Hamiltonian
$H=$ $-\frac{J}{2}\sum_{l=1}^{L-1}(\sigma^{x}_{l}\sigma_{l+1}^{x}+\sigma^{y}_{l}\sigma_{l+1}^{y}+\Delta\sigma^{z}_{l}\sigma_{l+1}^{z})$
$+h\sum_{l=1}^{L}l\sigma^{+}_{l}\sigma^{-}_{l}\leavevmode\nobreak\ ,$ (6)
where $\Delta$ represents the strength of the interaction between particles and $\sigma_{l}^{x,y,z}$ are the Pauli matrices acting on site $l$. In the single excitation subspace and with $\Delta{=}0$, the Hamiltonian in Eq. (4) reduces to the single-particle case in Eq. (4). We consider the time evolution of a system initially prepared in a Neel state, i.e. $\mathinner{|{\Psi(0)}\rangle}=\mathinner{|{1,0,\cdots,1,0}\rangle}$ for even $L$ and $\mathinner{|{\Psi(0)}\rangle}=\mathinner{|{1,0,\cdots,0,1}\rangle}$ for odd $L$. More details about the time evolution and complex pattern of many-body Bloch oscillations are available in Appendix B. We first put our focus on $\Delta{=}0$ (non-interacting case) and $\Delta{=}1$ (isotropic Heisenberg Hamiltonian). The results for the more general case of $10{<}\Delta{<}$ 1 are presented and discussed in the next section. In Figs. 3(a) and 3(b), we plot the long time limit of $\mathcal{F}_{Q}/t^{2}$ as a function of $h/J$ for various system sizes $L$ for $\Delta{=}0$ and $\Delta{=}1$, respectively. In both interacting and non-interacting regimes, the normalized QFI peaks at the transition point $h_{c}$ and it shows algebraic decay of the form $\mathcal{F}_{Q}/t^{2}\sim h^{-2}$ in the localized phase, which is similar to the single excitation case. However, unlike the single excitation scenario, even in the localized phase the QFI shows size dependence. Furthermore, one can see that in the presence of interaction, the behavior of the QFI qualitatively changes as: (i) it shows fluctuations in the vicinity of the transition point; and (ii) its value becomes smaller than the non-interacting case (i.e. $\Delta=0$). To extract the scaling with system size, in Figs. 3(c) and 3(d) we plot the long-time value of $\mathcal{F}_{Q}/t^{2}$ at $h{=}h_{c}$ as a function of system size $L$ for $\Delta{=}0$ and $\Delta{=}1$, respectively. Interestingly, in both cases, the scaling suggests $\beta>2$, which indicates a super-Heisenberg scaling. Especially, in the non-interacting case $\Delta=0$, it takes the form of $\mathcal{F}_{Q}(h_{c})\sim t^{2}L^{3}$. The difference of $\beta>2$ here, compared to the single-particle result of $\beta{=}2$ at the transition point, is related to the number of excitations that will be further explained in the next section.